One mole of oxygen at STP is adiabatically compressed to 5 atm. Calculate the final temperature. Take y = 1.4 and R= 8.31 J mol^-1 K^-1
1
Expert's answer
2019-05-21T08:43:59-0400

The relation between pressure and volume in an adiabatical process is

"P V^{\gamma} = \mathrm{const}"

Using

"PV = \nu RT"

we obtain

"P \bigg(\frac{T}{P}\bigg)^\gamma = \frac{T^{\gamma}}{P^{\gamma-1}}= \mathrm{const}"

or

"\frac{T_{\mathrm{final}}} {T_{\mathrm{initial}}}=\bigg(\frac{P_{\mathrm{final}}}{P_{\mathrm{initial}}}\bigg)^{\frac{\gamma-1}{\gamma}}"

so

"T_{\mathrm{final}} = T_{\mathrm{initial}} 5^{0.4/1.4} = 1.58 T_{\mathrm{initial}} = 432\, \mathrm{K}"